# A Large Recurrent Action Model: xLSTM Enables Fast Inference for Robotics Tasks

### v2

@@ D.5 Inference Time Comparisons
We empirically examine the difference in inference speed between of our models. Similar to De et al. (2024), we report both latency and throughput. For real-time applications, latency is the more important dimension, and therefore we focus our analysis on latency.
@@ D.5.1 Latency
In Figures 17 and 18, we report the latencies for DT and xLSTM with the same number of layer blocks as DT, and twice the number of layers blocks as DT, respectively. We conduct our comparison for two different batch sizes and across varying sequence lengths.
@@
Figure 17: Latency. We report latency with (a) batch size of 1 and (b) batch size of 16 for DT and xLSTM with 206M parameters. For xLSTM we use the same number of layer blocks as DT and a higher hidden dimension to match parameters.
@@ D.5.2 Throughput
In Figures 19 and 20, we similarly report the attained throughput for DT and xLSTM with the same number of layer blocks as DT, and twice the number of layers blocks as DT, respectively. We conduct our comparison for two fixed context lengths and varying batch sizes.
@@
Figure 19: Throughput. We report throughput with (a) context size of 800, and (b) context size of 1600 timesteps for DT and xLSTM with 206M parameters. For xLSTM we use the same number of layer blocks as DT and a higher hidden dimension to match parameters.
@@
D.5.3 xLSTM Kernel Comparisons
We leverage custom kernels for xLSTM to conduct our inference-speed comparisons. In particular, we compare 4 variants: recurrent-style inference with and without kernel acceleration, and chunkwise inference with and without kernel acceleration. In our experiments, every timestep contains 3 individual tokens. Consequently, regular recurrent-style inference requires iterating over the token sequence of length 3 in a loop given the hidden state of the previous timestep. This requires 3 forward passes. In contrast, the chunkwise implementation operates on chunks of timesteps given a hidden state. Consequently, this only requires a single forward pass. In Figure 21, we illustrate the impact of kernel acceleration. We find that our chunkwise kernels result in considerably lower latencies. Interestingly, we find that for $B=1$, our chunkwise implementation without kernel acceleration is faster than the recurrent-style inference with kernel acceleration. However, as the batch size increases, this trend reverses. This highlights the importance of kernel acceleration for efficient inference.
@@
Figure 21: Impact of kernel acceleration. We report latency with (a) batch size of 1 and (b) batch size of 32 for DT and xLSTM with 206M parameters. For xLSTM we use the same number of layer blocks as DT and a higher hidden dimension to match parameters.
@@ E.1.1 DT on Meta-World
We found that removing actions from the context results in better performance across backbones. In Figure 22, we report the learning curves over 200K updates for DT with varying context lengths on Meta-World, both with and without actions in the context. While context lengths beyond 1 hurt performance when training with actions, the reverse is true when training without actions. This is in contrast to recent works, which did not benefit from longer contexts (Octo Model Team et al., 2024). However, while removing actions improves performance on Meta-World, it does not affect performance on discrete control. On Meta-World, we observed that the models become overly confident (high action logits), which is problematic if poor initial actions are produced. We assume this is because in robotics actions change smoothly and by observing previous actions the agent learns shortcuts. A similar issue has been identified by Wen et al. (2020), and termed the copycat problem, because the agent is incentivized to copy previous actions. Our solution is to remove actions from the input sequence. This prevents the agent from learning shortcuts and alleviates the copycat problem.
@@
To further investigate the effect of removing actions from the context, we repeat this ablation on the full 432 tasks and 6 domains at the 206M model scale. In Figure 23, we report the learning curves for a DT with varying sequence lengths trained (a) with and (b) without actions in the agent’s context. Similar to the single-domain study on Meta-World with smaller models, we find that providing a longer context does not improve performance, resulting in a normalized score of around 0.3 across domains. In contrast, without action in the context, we observe a consistent improvement in the evaluation performance as the sequence length increases. In fact, the normalized score increase from around 0.3 with $C=1$ to 0.7 with $C=50$. For computational reasons, we only report one seed per sequence length in this experiment, but we believe that the overall trends are clear.
@@
To better understand on which domains the longer context benefits or hurts our agents, we also present the normalized score per domain in Figure 24. Without actions in the context, we find that longer context consistently benefits the performance across domains. With actions in the context we observe that on Meta-World and DMControl, the performance deteriorates for $C>1$. In contrast, on the discrete control domains Atari and Procgen, but also on the continuous continous control domain Composuite, performance tends to improve with $C>1$. This suggests that the copycat problem is particularly present on Meta-World and DMControl. However, note that the final performances on Atari, Procgen and Mimicgen are considerably worse when actions are present in the context compared to when they are not.
@@
To validate that modern recurrent backbones also benefit from training with longer sequence lengths, we repeat the same ablation as presented in Appendix E.1.2 using xLSTM [1:0]. We report the learning curves validation perplexities and evaluation performance across all 432 tasks for varying context lengths in Figure 25. Note that the validation perplexity curves in Figure 25a, start at step 50K for readability. Again, we observe considerable improvements in the validation perplexities and in the normalized scores (0.4 for $C=1$ to 0.8 for $C=50$) as the context length increases.
@@
Figure 25: Ablation on the effect of for varying the context length $C$ for xLSTM. We report (a) validation perplexity and (b) evaluation performance across the 432 training tasks for xLSTM [1:0]. Without actions in the context, the performance of DT improves with increasing $C$.
@@
Figure 26: Ablation on the effect of for varying the context length $C$ for xLSTM. We show the normalized scores per domain for all context lengths.
@@
To better understand whether our findings transfer to the behavior cloning setting, we conduct an ablation study in which we exclude the RTG tokens and the reward tokens from the sequence representation. This means that the sequence consists of state and reward tokens, or state-tokens only. In Figures 27 and 27, we report the (a) validation perplexities and (b) evaluation performance on the 432 task for the four considered backbones when removing RTG or RTG and reward, respectively. We retain the same training settings and datasets as reported in Appendix C (200K updates, evaluation after every 50K steps). We observe similar learning dynamics as for the 206M models that include RTG/reward tokens in the sequence representation (see Figure 2 and Figure 11). Consequently, we conclude that the same performance trends holds for training the considered backbones with and without RTG/reward condition. Note, that the final performances are lower compared to the models that include the RTG condition and that can be conditioned on a high return at inference time.
@@
Throughout our experiments, we compare two xLSTM variants: xLSTM [7:1] and xLSTM [1:0]. The bracket notation was introduced by (Beck et al., 2024), and denotes the ratio of mLSTM to sLSTM blocks. For example, xLSTM [7:1] contains 1 sLSTM block for every 7 mLSTM blocks. As described in Appendix C, we aim to maintain the same ration as proposed by Beck et al. (2024). While mLSTM blocks are fully parallelizable, sLSTM blocks are not. However, sLSTM preserves the non-diagonalized recurrent matrix to enable state-tracking (Merrill et al., 2024). As such, sLSTM can be attractive for tasks that require state-tracking (see Figure 4 in Beck et al. (2024)).
We first conduct an ablation study on the effect of the mLSTM-to-sLSTM ratio on the evaluation performance across all 432 tasks. For this experiment, we use the 16M parameter model that contains 8 xLSTM blocks in total. Consequently, we compare the following ratios [1:0] (only mLSTM), [0:1] (only sLSTM), [1:1], [1:3], [7:1]. In addition, we investigate the placement of sLSTMs across all 8 blocks. To indicate the placement, we use @ followed by the layer index (starting at 0). For example, [3:1] @ 1,3 indicates that the second and fourth layer are sLSTMs. In Figure 29 we report the validation perplexities and evaluation performance for different ratios and layer placements across the 432 tasks. For computational reasons, we conduct this experiment with only 1 seed per ratio. We find that at the 16M parameter scale, xLSTM [1:0] on average outperforms the variants that leverage sLSTM blocks. This indicates that these domains do not strongly benefit from the state tracking abilities of sLSTM.
@@
Next, conduct the same analysis on Dark-Room $10\times 10$ ICL environment as used in Appendix D.4. Unlike most of the 432 tasks used in our main experiments, Dark-Room exhibits a partially-observable observation space and sparse rewards. Consequently, Dark-Room is more likely to require state tracking abilities. In fact, we already observed better performance for xLSTM [7:1] than for xLSTM [1:0] in Appendix 16. In Figure 30, we report the ICL curves for the 80 train tasks and 20 hold-out tasks. We observe that xLSTM variants that contain sLSTM blocks at lower-level positions, such as [7:1] @ 1 and [3:1] @ 1,3 outperform xLSTM [1:0]. In contrast, xLSTM variants that contain sLSTM blocks at deeper-level positions, such as [0:1] and 3:1 @ 5,7, perform poorly. This is similar to findings by Beck et al. (2024) who also place sLSTM layers at lower-level positions.
@@ E.5 Effect of reducing number of layers in xLSTM
In Figure 32, report the validation perplexities and evaluation performance for the regular xLSTM with twice the number of layer blocks as DT, and an xLSTM with half the number of blocks. Reducing the number of layer blocks results in slight decrease in performance on both metrics. However, xLSTM still outperforms the Transformer baseline (see Figure 2).
@@
Figure 32: Ablation on the effect of reducing the number of layer blocks in xLSTM. We show the (a) validation perplexity and (b) evaluation performance on the training tasks for the layer regular and layer-matched matched xLSTM models. Reducing the number of layer blocks in xLSTM results in a slight performance decrease.
@@ Appendix F Embedding Space Analysis
In Figure 5, we analyze the representations learned by our models using UMAP (McInnes et al., 2018). Here, we explain the clustering procedure in more detail. For every task, we sample 32 sub-trajectories containing 50 timesteps (150 tokens) and encode them using our sequence models. Then, we extract the hidden states at the last layer of our model and aggregate them via mean pooling. We cluster all vectors using default hyperparameters of UMAP into a two-dimensional space. Finally, we color the resulting points by their domain. Generally, we find that tasks from the same domain cluster together.

### v3

@@ D.5 Inference Time Comparisons
We empirically examine the difference in inference speed between of our models. Similar to De et al. (2024), we report both latency and throughput. For real-time applications, latency is the more important dimension, and therefore, we focus our analysis on latency.
@@ D.5.1 Latency
In Figures 17 and 18, we report the latencies for DT and xLSTM with the same number of layer blocks as DT, and twice the number of layer blocks as DT, respectively. We conduct our comparison for two different batch sizes and across varying sequence lengths.
@@
Figure 17: Latency. We report latency with (a) batch size of 1 and (b) batch size of 16 for DT and xLSTM with 206M parameters. For xLSTM, we use the same number of layer blocks as DT and a higher hidden dimension to match parameters.
@@ D.5.2 Throughput
In Figures 19 and 20, we similarly report the attained throughput for DT and xLSTM with the same number of layer blocks as DT, and twice the number of layer blocks as DT, respectively. We conduct our comparison for two fixed context lengths and varying batch sizes.
@@
Figure 19: Throughput. We report throughput with (a) context size of 800, and (b) context size of 1600 timesteps for DT and xLSTM with 206M parameters. For xLSTM, we use the same number of layer blocks as DT and a higher hidden dimension to match parameters.
@@
D.5.3 xLSTM: Kernel Comparisons
We leverage custom kernels for xLSTM to conduct our inference-speed comparisons. In particular, we compare 4 variants: recurrent-style inference with and without kernel acceleration, and chunkwise inference with and without kernel acceleration. In our experiments, every timestep contains 3 individual tokens. Consequently, regular recurrent-style inference requires iterating over the token sequence of length 3 in a loop, given the hidden state of the previous timestep. This requires 3 forward passes. In contrast, the chunkwise implementation operates on chunks of timesteps given a hidden state. Consequently, this only requires a single forward pass. In Figure 21, we illustrate the impact of kernel acceleration. We find that our chunkwise kernels result in considerably lower latencies. Interestingly, we find that for $B=1$, our chunkwise implementation without kernel acceleration is faster than the recurrent-style inference with kernel acceleration. However, as the batch size increases, this trend reverses. This highlights the importance of kernel acceleration for efficient inference.
@@
Figure 21: Impact of kernel acceleration. We report latency with (a) batch size of 1 and (b) batch size of 32 for DT and xLSTM with 206M parameters. For xLSTM, we use the same number of layer blocks as DT and a higher hidden dimension to match parameters.
D.5.4 xLSTM: Impact of Head Dimension
In our experiments, we found that choosing the appropriate head dimension is critical to enable high throughput for xLSTM. Therefore, we conduct an inference ablation with xLSTM 206M in which we vary the number of heads between 4 and 32, while keeping the total hidden dimension constant, resulting in different head dimensions. We find that throughput increases considerably when increasing the number of heads (see Figure 22). For 4 heads, and therefore the highest head dimension, the total throughput saturates at batch size 96. In contrast, when increasing the number of heads to 32 (i.e., decreasing the head dimension), the total throughput continues to increase. This is because a higher head dimension incurs more FLOPS.
Figure 22: Throughput comparison for xLSTM 206M with varying numbers of heads but fixed total hidden size. By default, we used 4 heads for our experiments. Increasing the number of heads results in higher throughput.
@@ E.1.1 DT on Meta-World
We found that removing actions from the context results in better performance across backbones. In Figure 23, we report the learning curves over 200K updates for DT with varying context lengths on Meta-World, both with and without actions in the context. While context lengths beyond 1 hurt performance when training with actions, the reverse is true when training without actions. This is in contrast to recent works, which did not benefit from longer contexts (Octo Model Team et al., 2024). However, while removing actions improves performance on Meta-World, it does not affect performance on discrete control. On Meta-World, we observed that the models become overly confident (high action logits), which is problematic if poor initial actions are produced. We assume this is because in robotics, actions change smoothly, and by observing previous actions, the agent learns shortcuts. A similar issue has been identified by Wen et al. (2020) and termed the copycat problem, because the agent is incentivized to copy previous actions. Our solution is to remove actions from the input sequence. This prevents the agent from learning shortcuts and alleviates the copycat problem.
@@
To further investigate the effect of removing actions from the context, we repeat this ablation on the full 432 tasks and 6 domains at the 206M model scale. In Figure 24, we report the learning curves for a DT with varying sequence lengths trained (a) with and (b) without actions in the agent’s context. Similar to the single-domain study on Meta-World with smaller models, we find that providing a longer context does not improve performance, resulting in a normalized score of around 0.3 across domains. In contrast, without action in the context, we observe a consistent improvement in the evaluation performance as the sequence length increases. In fact, the normalized score increases from around 0.3 with $C=1$ to 0.7 with $C=50$. For computational reasons, we only report one seed per sequence length in this experiment, but we believe that the overall trends are clear.
@@
To better understand on which domains the longer context benefits or hurts our agents, we also present the normalized score per domain in Figure 25. Without actions in the context, we find that longer context consistently benefits the performance across domains. With actions in the context, we observe that on Meta-World and DMControl, the performance deteriorates for $C>1$. In contrast, on the discrete control domains Atari and Procgen, but also on the continuous control domain Composuite, performance tends to improve with $C>1$. This suggests that the copycat problem is particularly present on Meta-World and DMControl. However, note that the final performances on Atari, Procgen, and Mimicgen are considerably worse when actions are present in the context compared to when they are not.
@@
To validate that modern recurrent backbones also benefit from training with longer sequence lengths, we repeat the same ablation as presented in Appendix E.1.2 using xLSTM [1:0]. We report the learning curves, validation perplexities, and evaluation performance across all 432 tasks for varying context lengths in Figure 26. Note that the validation perplexity curves in Figure 26a, start at step 50K for readability. Again, we observe considerable improvements in the validation perplexities and the normalized scores (0.4 for $C=1$ to 0.8 for $C=50$) as the context length increases.
@@
Figure 26: Ablation on the effect of varying the context length $C$ for xLSTM. We report (a) validation perplexity and (b) evaluation performance across the 432 training tasks for xLSTM [1:0]. Without actions in the context, the performance of DT improves with increasing $C$.
@@
Figure 27: Ablation on the effect of varying the context length $C$ for xLSTM. We show the normalized scores per domain for all context lengths.
@@
To better understand whether our findings transfer to the behavior cloning setting, we conduct an ablation study in which we exclude the RTG tokens and the reward tokens from the sequence representation. This means that the sequence consists of state and reward tokens, or state-tokens only. In Figures 28 and 28, we report the (a) validation perplexities and (b) evaluation performance on the 432 task for the four considered backbones when removing RTG or RTG and reward, respectively. We retain the same training settings and datasets as reported in Appendix C (200K updates, evaluation after every 50K steps). We observe similar learning dynamics as for the 206M models that include RTG/reward tokens in the sequence representation (see Figure 2 and Figure 11). Consequently, we conclude that the same performance trends hold for training the considered backbones with and without RTG/reward condition. Note that the final performances are lower compared to the models that include the RTG condition, and that can be conditioned on a high return at inference time.
@@
Throughout our experiments, we compare two xLSTM variants: xLSTM [7:1] and xLSTM [1:0]. The bracket notation was introduced by (Beck et al., 2024) and denotes the ratio of mLSTM to sLSTM blocks. For example, xLSTM [7:1] contains 1 sLSTM block for every 7 mLSTM blocks. As described in Appendix C, we aim to maintain the same ratio as proposed by Beck et al. (2024). While mLSTM blocks are fully parallelizable, sLSTM blocks are not. However, sLSTM preserves the non-diagonalized recurrent matrix to enable state-tracking (Merrill et al., 2024). As such, sLSTM can be attractive for tasks that require state-tracking (see Figure 4 in Beck et al. (2024)).
We first conduct an ablation study on the effect of the mLSTM-to-sLSTM ratio on the evaluation performance across all 432 tasks. For this experiment, we use the 16M parameter model that contains 8 xLSTM blocks in total. Consequently, we compare the following ratios [1:0] (only mLSTM), [0:1] (only sLSTM), [1:1], [1:3], [7:1]. In addition, we investigate the placement of sLSTMs across all 8 blocks. To indicate the placement, we use @ followed by the layer index (starting at 0). For example, [3:1] @ 1,3 indicates that the second and fourth layers are sLSTMs. In Figure 30, we report the validation perplexities and evaluation performance for different ratios and layer placements across the 432 tasks. For computational reasons, we conduct this experiment with only 1 seed per ratio. We find that at the 16M parameter scale, xLSTM [1:0] on average outperforms the variants that leverage sLSTM blocks. This indicates that these domains do not strongly benefit from the state tracking abilities of sLSTM.
@@
Next, conduct the same analysis on Dark-Room $10\times 10$ ICL environment as used in Appendix D.4. Unlike most of the 432 tasks used in our main experiments, Dark-Room exhibits a partially observable observation space and sparse rewards. Consequently, Dark-Room is more likely to require state tracking abilities. In fact, we already observed better performance for xLSTM [7:1] than for xLSTM [1:0] in Appendix 16. In Figure 31, we report the ICL curves for the 80 train tasks and 20 hold-out tasks. We observe that xLSTM variants that contain sLSTM blocks at lower-level positions, such as [7:1] @ 1 and [3:1] @ 1,3 outperform xLSTM [1:0]. In contrast, xLSTM variants that contain sLSTM blocks at deeper-level positions, such as [0:1] and 3:1 @ 5,7, perform poorly. This is similar to findings by Beck et al. (2024) who also place sLSTM layers at lower-level positions.
@@ E.5 Effect of reducing number of layers in xLSTM
In Figure 33, report the validation perplexities and evaluation performance for the regular xLSTM with twice the number of layer blocks as DT, and an xLSTM with half the number of blocks. Reducing the number of layer blocks results in a slight decrease in performance on both metrics. However, xLSTM still outperforms the Transformer baseline (see Figure 2).
@@
Figure 33: Ablation on the effect of reducing the number of layer blocks in xLSTM. We show the (a) validation perplexity and (b) evaluation performance on the training tasks for the layer regular and layer-matched xLSTM models. Reducing the number of layer blocks in xLSTM results in a slight performance decrease.
@@ Appendix F Embedding Space Analysis
In Figure 5, we analyze the representations learned by our models using UMAP (McInnes et al., 2018). Here, we explain the clustering procedure in more detail. For every task, we sample 32 sub-trajectories containing 50 timesteps (150 tokens) and encode them using our sequence models. Then, we extract the hidden states at the last layer of our model and aggregate them via mean pooling. We cluster all vectors using the default hyperparameters of UMAP into a two-dimensional space. Finally, we color the resulting points by their domain.
The purpose of this analysis is to examine how the models organize their representations of different environments. In general, tasks within the same domain tend to share similar input characteristics, such as visual inputs (e.g., image frames), possible actions to perform, and reward structures. Therefore, they are more likely to be “grouped” together in the embedding space. For example, when embeddings of Atari games are closer to each other than to Procgen games, it indicates that Atari games share more similar underlying dynamics or input structures compared to Procgen. We indeed find that tasks from the same domain cluster together. A more refined and better-separated embedding space may result in better final performance, potentially because it facilitates task identification at inference time. This may, however, be specific to the mixture of training tasks at hand. Therefore, we believe that studying the learned embedding spaces of multi-task agents in a wide range of environments is interesting for future work.
Analogous to Figure 5 for DT and xLSTM, we show the UMAP clustering for Mamba 16M in Figure 34. In comparison to DT, Mamba exhibits a slightly stronger grouping of the embedding space.
(a) DT
(b) Mamba
(c) xLSTM
Figure 34: UMAP clustering of hidden states for 432 tasks produced by (a) DT, (b) Mamba, and (c) xLSTM with 16M parameters, colored by domain. We again depict the embedding spaces for DT and xLSTM from Figure 5 for better readability.